if G is centroid of triangleABC and if G1 be the centroid of another trianglePQR show that AP(vector)+BQ(vector)+CR(vector)=3(G)(G1)

Grade:12

2 Answers

bharat bajaj IIT Delhi
10 years ago
The capital letters denote the position vectors of the points:
G = A + B + C/3
G1 = P + Q + R/3
3(G-G1) = A-P + B - Q + C - R
3GG1 = AP + BQ + CR
Thanks
Bharat Bajaj
bharat bajaj IIT Delhi
10 years ago
The capital letters denote the position vectors of the points:
G = A + B + C/3
G1 = P + Q + R/3
3(G-G1) = A-P + B - Q + C - R
3GG1 = AP + BQ + CR

Thanks
Bharat Bajaj
askiitians faculty
IIT Delhi